OpenStudy (anonymous):

find the sum of the given polynomials. 2mn + 4n 2 and 5m 2 - 7mn A) 9m2 + 9mn - 4n2 B) 5m2 - 14mn - 4n2 C) 5m2 - 5mn + 4n2 D) 5m2 + 5mn - 4n2

OpenStudy (anonymous):

do you need help

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok just give me a moment

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so first you have to add that right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after that you will get 5mn+4n2+5m2

OpenStudy (anonymous):

is that one of your answers

OpenStudy (anonymous):

c.)

OpenStudy (anonymous):

No @victoriajustice

OpenStudy (anonymous):

somebody just said c

OpenStudy (anonymous):

=c.)

OpenStudy (anonymous):

me jordan did

OpenStudy (anonymous):

Ok @my.baby

OpenStudy (anonymous):

victoria will you be my fan

OpenStudy (anonymous):

i see why I added to much

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

i am your fan and gave you a medal

OpenStudy (anonymous):

what number is your smartscore

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i donteven know what that is

OpenStudy (anonymous):

you mind if you help me

OpenStudy (anonymous):

the number at the top of the screen

OpenStudy (anonymous):

i don't mind

OpenStudy (anonymous):

24

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

should i ask you the question now

OpenStudy (anonymous):

how do you give medals

OpenStudy (anonymous):

yyaass

OpenStudy (anonymous):

A teacher wrote expressions on cards and put them in a bag. Thomas and Kelly drew cards from the bag and added the expressions. Thomas drew the expressions 3x + 7 and 2x – 3. Kelly drew the expressions −6x + 3 and 14x + 2. How much greater was Kelly’s sum than Thomas’s sum? x +

OpenStudy (anonymous):

on the side of my name it should say best response in gray

OpenStudy (anonymous):

i think i have to subtract

OpenStudy (anonymous):

yea i think so too

OpenStudy (anonymous):

i got 20x-5

OpenStudy (anonymous):

i dont think that is right

OpenStudy (anonymous):

are you still here

OpenStudy (anonymous):

ccccccccc

OpenStudy (anonymous):

c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

this is not an abc question it is type it in

OpenStudy (anonymous):

whats that for

OpenStudy (anonymous):

ok let me look again

OpenStudy (anonymous):

it says type it in i wish i did have abc and d i am sorry

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

-x+-

OpenStudy (anonymous):

we have to figure that out

OpenStudy (anonymous):

its going to take me a little time to figure it out but i am good at algerbra

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you still here

OpenStudy (anonymous):

i think its addition

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

it is 13x+6

OpenStudy (anonymous):

+

OpenStudy (anonymous):

yes i added

OpenStudy (anonymous):

lets do another one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = x – 3 ft s2 = x + 2 ft s3 = x + 4 ft P = x + ft

OpenStudy (anonymous):

6

OpenStudy (anonymous):

on both

OpenStudy (anonymous):

6x+6ft

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have to go

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

see you later i guess

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

bye